OpenStudy (anonymous):

Three-sevenths of a number is 21. Find the number. the number is 9 the number is 27 the number is 49

OpenStudy (aric200):

the number is 49

OpenStudy (aric200):

x=21*7/3 x=147/3 x=49

OpenStudy (anonymous):

yes 49 y ese wey me copio !!!

OpenStudy (anonymous):

ok prima thanks aniyways

OpenStudy (aric200):

All you did was copy my answer and you got the Medal...

OpenStudy (anonymous):

anww chilax its the sameso keep it cool

OpenStudy (aric200):

-_-

OpenStudy (anonymous):

:P